A man rowing at the rate of 5 km an hour in still water takes thrice as much time in going 40 km up the river as in going 40 km down . Find the rate at which the river flows ? ( Sir pls give me an quick answer ) (thanks )
Answer by checkley71(8403)

EQUATIONS ARE UPSTREAM (5-C) C IS THE CURRENT AND DOWN STREAM (5+C) AND IT TAKES TWICE AS LONG TO GO UP STREAM THAN DOWN STREAM THEREFORE
2(5-C)=(5+C) OR 10-2C=5+C OR -3C=-5 OR C=5/3KM AN HOUR
PROOF 2(5-5/3)=5+5/3 OR 10-10/3=5+5/3 OR 10-3 1/3=5+5/3 OR 6 2/3 =6 2/3